Calculate the force of gravity between a newborn baby of mass 3.3 kg and the obstetrician of mass 100 kg , who is 0.50 m from the baby.
1
Expert's answer
2019-09-30T10:08:42-0400
"F=\frac{GmM}{r^2}"

"F=\frac{(6.67\cdot 10^{-11})(3.3)(100)}{0.5^2}=8.8\cdot 10^{-8}N"